OpenStudy (petlovers123):

A rectangular prism is completely packed with 200 cubes of edge length fraction 1/5 inch, without any gap or overlap. Which of these best describes the volume of this rectangular prism? 1 unit cube and 15 smaller cubes of volume fraction 1/125 cubic inch each 1 unit cube and 75 smaller cubes of volume fraction 1/125 cubic inch each 7 unit cubes and 25 smaller cubes of volume fraction 1/125 cubic inch each 7 unit cubes and 125 smaller cubes of volume fraction 1/125 cubic inch each

OpenStudy (phi):

do you know the formula for the volume of a cube ?

OpenStudy (petlovers123):

nope, Thats all the question brings

OpenStudy (phi):

yes, I see that. But you are supposed to bring the formula for the volume of a cube in other words, they expect you to know that. can you look it up ?

OpenStudy (petlovers123):

Volume of a cube = side times side times side. Since each side of a square is the same, it can simply be the length of one side cubed. If a square has one side of 4 inches, the volume would be 4 inches times 4 inches times 4 inches, or 64 cubic inches.

OpenStudy (phi):

ok, using that , what is the volume of one of the tiny cubes they mention ?

OpenStudy (petlovers123):

16

OpenStudy (phi):

cubes of edge length fraction 1/5 inch they are saying each side is 1/5 inch long.

OpenStudy (phi):

Volume of a cube = side times side times side

OpenStudy (petlovers123):

oh ok

OpenStudy (phi):

any idea what the volume of that cube is ?

OpenStudy (petlovers123):

is it 0.8 inches

OpenStudy (phi):

what is the length of one side ?

OpenStudy (petlovers123):

1/5 inch long

OpenStudy (phi):

and what is side * side * side when side is 1/5 ?

OpenStudy (petlovers123):

so i multiply 1/5 x 3 or 4

OpenStudy (phi):

you are doing \[ \frac{1}{5} \cdot \frac{1}{5}\cdot \frac{1}{5} \] first do \[ \frac{1}{5} \cdot \frac{1}{5} \]

OpenStudy (phi):

when multiplying fractions, multiply top times top and bottom times bottom

OpenStudy (petlovers123):

i multiplied 1/5 x 1/5 and I got 0.04

OpenStudy (phi):

I would keep it as a fraction, because the answer choices are fractions what is the answer as a fraction? (people learn how to do this in their head) the top is 1*1 the bottom is 5*5

OpenStudy (petlovers123):

1/25

OpenStudy (petlovers123):

1/125

OpenStudy (phi):

yes 1/5*1/5 is 1/25 and 1/25* 1/5 is 1/125 that is the volume of one cube any idea how to find the volume of 200 of these cubes ? (I would multiply 200*1/125)

OpenStudy (petlovers123):

1.6

OpenStudy (phi):

ok, but it's still hard to pick out the correct choice I think they want you to do \[ 200 \cdot \frac{1}{125}= \frac{200}{125} \] 125 divides into 200 1 with a remainder. any idea what the remainder is ?

OpenStudy (petlovers123):

i dont get the 200 1 part

OpenStudy (phi):

in other words, 125 goes into 200 once but there is some amount left over

OpenStudy (phi):

Have you ever seen